OpenStudy (anonymous):

I Honestly Need Help Find The Equestion M= 4; (-1 1over 2 )

OpenStudy (anonymous):

Thats What I Said

OpenStudy (katrinakaif):

Use the equation (y-y1) = m (x-x1) m = 4 while.. (-1, 1/2) = (x1, y1)

OpenStudy (anonymous):

what would be the answer